OpenStudy (anonymous):

int ((3)/(4-7x)) dx Work would be great, please and ty

OpenStudy (amistre64):

your missing a -7

OpenStudy (amistre64):

\[-\frac{3}{7}\int\frac{-7}{4-7x}dx \]

OpenStudy (anonymous):

What do you mean i was missing a -7?

OpenStudy (amistre64):

i mean that if you had a -7 you could simply ln it up

OpenStudy (amistre64):

we can always apply a useful form of "1" to borrow a constant from without changing the value of anything

OpenStudy (amistre64):

\[\frac{-7}{-7}\int\frac{3}{4-7x}dx\to \ ` -\frac{3}{7}\int\frac{-7}{4-7x}dx\]

OpenStudy (dape):

Move the 3 outside of the integral (it's constant) and do the substitution \[ u=4-7x \\ du = -7dx \] or \[ dx = \frac{du}{-7} \] We now get: \[ 3\int{\frac{1}{u}\frac{du}{-7}} = -\frac{3}{7}\int{\frac{du}{u}} = -\frac{3}{7}\ln{u} \] By putting our substitution back we get: \[ \int{\frac{3}{4-7x}dx} = -\frac{3}{7}\ln|4-7x| \]

OpenStudy (anonymous):

Great explanation, thank you!

OpenStudy (dape):

No problem. Also, if you didn't already know, the absolute value inside the ln is required because it isn't defined for negative values.